OpenStudy (anonymous):

medal please help me!selling an item for rs1800 at a discount of 10% a shopkeeper had a gain of rs200.had he sold the item without discount what would be the percentage of profit

OpenStudy (anonymous):

LLAMAZ

OpenStudy (anonymous):

means what

OpenStudy (anonymous):

LLAMAZ ARE THE ANSWER TO EVERYTHING OK

OpenStudy (anonymous):

1800*100/90=2000

OpenStudy (anonymous):

NO

OpenStudy (anonymous):

YOUR WRONG

OpenStudy (anonymous):

LLAMAZ ARE THE ANSWER

OpenStudy (anonymous):

I dont know internet slang @KraziiKai .i am from india

OpenStudy (anonymous):

cost price=1600 ,selling price =1980 then gain %? gain%=380/1600*100=23.75

OpenStudy (anonymous):

@tejasvir

OpenStudy (gorv):

welcome to os rosy :)

OpenStudy (anonymous):

hi gorv

OpenStudy (gorv):

heyy

OpenStudy (anonymous):

tell me the meaning of llamaz

OpenStudy (anonymous):

Is it a bad word

OpenStudy (anonymous):

means

OpenStudy (anonymous):

means mouth of animal why she is using here

OpenStudy (gorv):

i dont know

OpenStudy (kropot72):

Cost price = 1800 - 200 = rs1600 Selling price without discount is given by: \[\large Selling\ price\ without\ discount=\frac{1800}{1.0-0.1}=\frac{1800}{0.9}=rs2000\] Percentage profit is sold without discount is: \[\large \frac{2000-1600}{1600}\times\frac{100}{1}=25\ percent\]

OpenStudy (anonymous):

yes i made a mistake in selling price.